# Memory-Efficient Training with In-Place FFT Implementation

### v1

@@ 1 Introduction
Large-scale neural models have achieved remarkable success across a wide range of applications, such as natural language processing [1][2] and computer vision [3]. As model sizes increase, memory consumption has emerged as a significant challenge. Notably, model training memory cost is larger than the deployment cost, primarily due to the backpropagation process [4]. Thus, reducing memory usage has become a critical research study, especially for the training stage.
Numerous methods have been proposed to reduce memory usage during model training and deployment. Commonly used approaches such as model distillation [5], quantization [6], and pruning [7] mainly aim to reduce memory consumption by decreasing the number or precision of model parameters. In contrast, this work takes a novel approach at the arithmetic operator level, implementing in-place Fast Fourier Transform (FFT) operations for model training through the use of circulant-structured parameter matrices [8, 9, 10]. The FFT operator has been widely employed in various neural network architectures due to its efficiency in capturing global patterns and enabling structured transformations. For instance, Fourier-based fine-tuning methods such as FourierFT [11] and Block Circulant Adapter (BCA) [10] rely heavily on FFT to perform efficient parameter transformations in the frequency domain.
It should be noted that FFT-based operators typically involve both FFT and IFFT computations, which generate intermediate tensors that consume memory and use different data type since the result of FFT is complex number. Numerous studies have explored memory optimization strategies for FFT operators, particularly in high-performance programming libraries such as FFTW [12] and cuFFT [13]. These libraries support in-place computation, where input and output share memory space. However, they suffer from the inability to maintain the original (input) memory space and lack of support for bfloat16 data type that is widely used for modern neural models.
@@ 2 Related Works
Current automatic differentiation frameworks often discourage in-place operations due to their challenges in gradient computation during training. The key benefit of in-place computation lies in its ability to save memory by eliminating the need for storing intermediate tensors. For instance, combining batch normalization and activation into a single in-place operation has been shown to reduce memory usage by up to half [14]. In neural network based hardware placement tasks [15], in-place operations are found beneficial due to their inherent memory efficiency. Similarly, in-place operations are integrated into convolutional neural networks for anomaly detection [16], achieving notable memory savings. The in-place operation can also be extended to broader contexts. For example, in in-place distillation [17], a student model is distilled directly from the teacher model without additional memory allocation.
The FFT operator can be found across various neural network models, particularly in tasks involving the Fourier domain, which is common in computer vision. Some convolutional neural networks run entirely in the Fourier domain [18]. In beampattern synthesis [19], an IFFT operator is applied to the hidden representations generated by the neural network. Additionally, 2D FFT has proven beneficial for fine-tuning large language models [11]. Circulant-structured matrix-vector products can also be efficiently computed using 1D FFT and IFFT, facilitating neural network training [10]. Beyond these applications, FFT also plays a key role in spectral convolution [20] and in approximating global attention mechanisms [21].
From the perspective of automatic differentiation, each operator can be seen as a neural network layer, with or without trainable parameters. For example, the low-rank fine-tuning method for large language models [22] represents a full weight matrix using two low-rank matrices, which is equivalent to introducing two linear layers, thereby requiring storage for intermediate activations. Similarly, circulant-structured weight matrices [10] leverage FFT and IFFT to transform inputs and parameters. Although FFT operators themselves does not contain trainable parameters, they still require the preservation of intermediate results to support automatic differentiation and to handle complex-valued data types that is different from the real-valued inputs. Moreover, libraries such as FFTW [12] and cuFFT [13] demand pre-allocated memory buffers of size $N+2$ real numbers. This memory pre-allocation should be handled during the neural model loading phase, which complicates integration and limits the practicality of using these libraries in real-world neural network applications. Besides, they also does not support bfloat16 data type which is common in modern neural networks.
@@ 3.1 Standard FFT and rFFT
Theorem 1 (Conjugate Symmetry of Real FFT [23]).
@@ 3.2 In-place Transform via Butterfly Operation
The FFT can be efficiently computed using the Cooley–Tukey algorithm [24], which recursively decomposes given FFT into smaller FFTs. As illustrated in the Butterfly Operation Diagram section of Fig.1, a 16-point FFT is first recursively decomposed into smaller FFTs—two 8-point FFTs, then four 4-point FFTs, and finally eight 2-point FFTs. Following this hierarchical decomposition, the FFT is computed through a series of butterfly operations applied at each level of the recursion. At the core of the algorithm lies the butterfly operation, a fundamental computation that transforms a pair of complex values into two outputs using addition, subtraction, and multiplication by a twiddle factor.
@@ 3.3 FFT based Model Training via Circulant Matrix
The circulant matrix based neural network training can be converted to FFT and IFFT operations for acceleration [8]. Given a real-valued circulant weight matrix $\mathbf{C}\in\mathbb{R}^{N\times N}$ defined by its first column $\mathbf{c}\in\mathbb{R}^{N}$, the linear transformation $\mathbf{y}=\mathbf{C}\mathbf{x}$ can be equivalently computed in the frequency domain as follows:
@@
This structure not only accelerates computation but also simplifies the gradient computation during training. By leveraging the linearity and conjugate of the FFT results [10], gradients with respect to both input $\mathbf{x}$ and parameter $\mathbf{c}$ can be computed by:
@@
where $\overline{\texttt{FFT}(\cdot)}$ means taking the conjugate of the FFT results. There is also block circulant matrix based training that aims to fit non-square matrix [9], where the matrix is divided into blocks by partition size $p$.
@@ 5 Experiments
All our experiments compare three different FFT implementations: (1) fft: standard complex-valued FFT using torch.fft.fft/ifft from PyTorch [25]; (2) rfft: real-input FFT using torch.fft.rfft/irfft, exploiting Hermitian symmetry; (3) ours: custom CUDA-based real-domain FFT with in-place forward/backward implementation, reusing the input real-valued memory for intermediate result storage.
We also include two common baselines: (1) FF: updating all trainable parameters; (2) lora[22]: low-rank adaptation with parameter-efficient updates. For fair comparison, all runs use the same training configuration (batch size, optimizer, precision).
@@ 5.1 Memory Efficiency
To evaluate the memory efficiency of our proposed in-place training, we conduct experiments in two settings: (1) single-layer analysis: we perform training on a singular linear layer with different training methods, where circulant matrix based training [10] are accomplished with different FFT implementations; (2) full-model training: we apply the circulant fine-tuning approach [10] to RoBERTa-large and LLaMA2-7B and monitor memory usage throughout training. All circulant variants share the same number of trainable parameters, differing only in FFT backend.
@@ 5.1.2 Full Model Training
Setups. We conduct full-model experiments on both LLaMA2-7B and RoBERTa-large using an NVIDIA A100 GPU. For LLaMA2-7B, we use the GSM8K dataset with per_device_train_batch_size set to 2 and gradient_accumulation_steps set to 4. For RoBERTa-large, we use the MRPC dataset with a batch size of 32. These configurations follow the standard precision training setups for each task [10], and we retain them to reflect real-world GPU memory cost. We use stochastic gradient descent (SGD) as the optimizer in all experiments.
@@ 5.2.2 Model-Level Evaluation
Table 4: Runtime throughput and MRPC accuracy of different fine-tuning methods. Token-level throughput (Thr., in k tokens/sec) is measured on LLaMA-2-7B using the GSM8K dataset with one A800 GPU, while MRPC classification accuracy (Acc., %) is evaluated on RoBERTa-large. Accuracy results are reported from the Block-Circulant Adapter (BCA) work [10], where only limited configurations were provided, leading to some “N/A” entries. All lora experiments use rank $r=32$. For circulant-based methods, p denotes the block size.

### v2

@@ 1 Introduction
Large-scale neural models have achieved remarkable success across a wide range of applications, such as natural language processing [GPT-3][T5] and computer vision [ViT]. As model sizes increase, memory consumption has emerged as a significant challenge. Notably, model training memory cost is larger than the deployment cost, primarily due to the backpropagation process [lecun1988theoretical]. Thus, reducing memory usage has become a critical research study, especially for the training stage.
Numerous methods have been proposed to reduce memory usage during model training and deployment. Commonly used approaches such as model distillation [hinton2015distilling], quantization [courbariaux2015low], and pruning [DBLP:journals/corr/HanMD15] mainly aim to reduce memory consumption by decreasing the number or precision of model parameters. In contrast, this work takes a novel approach at the arithmetic operator level, implementing in-place Fast Fourier Transform (FFT) operations for model training through the use of circulant-structured parameter matrices [cheng2015exploration, ding2017circnn, circulant]. The FFT operator has been widely employed in various neural network architectures due to its efficiency in capturing global patterns and enabling structured transformations. For instance, Fourier-based fine-tuning methods such as FourierFT [fourierft] and Block Circulant Adapter (BCA) [circulant] rely heavily on FFT to perform efficient parameter transformations in the frequency domain.
It should be noted that FFT-based operators typically involve both FFT and IFFT computations, which generate intermediate tensors that consume memory and use different data type since the result of FFT is complex number. Numerous studies have explored memory optimization strategies for FFT operators, particularly in high-performance programming libraries such as FFTW [fftw] and cuFFT [cufft_layout]. These libraries support in-place computation, where input and output share memory space. However, they suffer from the inability to maintain the original (input) memory space and lack of support for bfloat16 data type that is widely used for modern neural models.
@@
The implementation of the proposed operators is being discussed for upstreaming into PyTorch. The corresponding discussion is available at: https://github.com/pytorch/pytorch/issues/171022.
@@ 2 Related Works
Current automatic differentiation frameworks often discourage in-place operations due to their challenges in gradient computation during training. The key benefit of in-place computation lies in its ability to save memory by eliminating the need for storing intermediate tensors. For instance, combining batch normalization and activation into a single in-place operation has been shown to reduce memory usage by up to half [bulo2018place]. In neural network based hardware placement tasks [liu2022xplace], in-place operations are found beneficial due to their inherent memory efficiency. Similarly, in-place operations are integrated into convolutional neural networks for anomaly detection [sun2023tinyad], achieving notable memory savings. The in-place operation can also be extended to broader contexts. For example, in in-place distillation [yu2019universally], a student model is distilled directly from the teacher model without additional memory allocation.
The FFT operator can be found across various neural network models, particularly in tasks involving the Fourier domain, which is common in computer vision. Some convolutional neural networks run entirely in the Fourier domain [pratt2017fcnn]. In beampattern synthesis [zhao2023efficient], an IFFT operator is applied to the hidden representations generated by the neural network. Additionally, 2D FFT has proven beneficial for fine-tuning large language models [fourierft]. Circulant-structured matrix-vector products can also be efficiently computed using 1D FFT and IFFT, facilitating neural network training [circulant]. Beyond these applications, FFT also plays a key role in spectral convolution [Spectral] and in approximating global attention mechanisms [Fnet].
From the perspective of automatic differentiation, each operator can be seen as a neural network layer, with or without trainable parameters. For example, the low-rank fine-tuning method for large language models [lora] represents a full weight matrix using two low-rank matrices, which is equivalent to introducing two linear layers, thereby requiring storage for intermediate activations. Similarly, circulant-structured weight matrices [circulant] leverage FFT and IFFT to transform inputs and parameters. Although FFT operators themselves does not contain trainable parameters, they still require the preservation of intermediate results to support automatic differentiation and to handle complex-valued data types that is different from the real-valued inputs. Moreover, libraries such as FFTW [fftw] and cuFFT [cufft_layout] demand pre-allocated memory buffers of size $N+2$ real numbers. This memory pre-allocation should be handled during the neural model loading phase, which complicates integration and limits the practicality of using these libraries in real-world neural network applications. Besides, they also does not support bfloat16 data type which is common in modern neural networks.
@@ 3.1 Standard FFT and rFFT
Theorem 1 (Conjugate Symmetry of Real FFT [conj_rDFT]).
@@ 3.2 In-place Transform via Butterfly Operation
The FFT can be efficiently computed using the Cooley–Tukey algorithm [cooley1965algorithm], which recursively decomposes given FFT into smaller FFTs. As illustrated in the Butterfly Operation Diagram section of Fig.1, a 16-point FFT is first recursively decomposed into smaller FFTs—two 8-point FFTs, then four 4-point FFTs, and finally eight 2-point FFTs. Following this hierarchical decomposition, the FFT is computed through a series of butterfly operations applied at each level of the recursion. At the core of the algorithm lies the butterfly operation, a fundamental computation that transforms a pair of complex values into two outputs using addition, subtraction, and multiplication by a twiddle factor.
@@ 3.3 FFT based Model Training via Circulant Matrix
The circulant matrix based neural network training can be converted to FFT and IFFT operations for acceleration [cheng2015exploration]. Given a real-valued circulant weight matrix $\mathbf{C}\in\mathbb{R}^{N\times N}$ defined by its first column $\mathbf{c}\in\mathbb{R}^{N}$, the linear transformation $\mathbf{y}=\mathbf{C}\mathbf{x}$ can be equivalently computed in the frequency domain as follows:
@@
This structure not only accelerates computation but also simplifies the gradient computation during training. By leveraging the linearity and conjugate of the FFT results [circulant], gradients with respect to both input $\mathbf{x}$ and parameter $\mathbf{c}$ can be computed by:
@@
where $\overline{\texttt{FFT}(\cdot)}$ means taking the conjugate of the FFT results. There is also block circulant matrix based training that aims to fit non-square matrix [ding2017circnn], where the matrix is divided into blocks by partition size $p$.
@@ 5 Experiments
All our experiments compare three different FFT implementations: (1) fft: standard complex-valued FFT using torch.fft.fft/ifft from PyTorch [paszke2017automatic]; (2) rfft: real-input FFT using torch.fft.rfft/irfft, exploiting Hermitian symmetry; (3) ours: custom CUDA-based real-domain FFT with in-place forward/backward implementation, reusing the input real-valued memory for intermediate result storage.
We also include two common baselines: (1) FF: updating all trainable parameters; (2) lora[lora]: low-rank adaptation with parameter-efficient updates. For fair comparison, all runs use the same training configuration (batch size, optimizer, precision).
@@ 5.1 Memory Efficiency
To evaluate the memory efficiency of our proposed in-place training, we conduct experiments in two settings: (1) single-layer analysis: we perform training on a singular linear layer with different training methods, where circulant matrix based training [circulant] are accomplished with different FFT implementations; (2) full-model training: we apply the circulant fine-tuning approach [circulant] to RoBERTa-large and LLaMA2-7B and monitor memory usage throughout training. All circulant variants share the same number of trainable parameters, differing only in FFT backend.
@@ 5.1.2 Full Model Training
Setups. We conduct full-model experiments on both LLaMA2-7B and RoBERTa-large using an NVIDIA A100 GPU. For LLaMA2-7B, we use the GSM8K dataset with per_device_train_batch_size set to 2 and gradient_accumulation_steps set to 4. For RoBERTa-large, we use the MRPC dataset with a batch size of 32. These configurations follow the standard precision training setups for each task [circulant], and we retain them to reflect real-world GPU memory cost. We use stochastic gradient descent (SGD) as the optimizer in all experiments.
@@ 5.2.2 Model-Level Evaluation
Table 4: Runtime throughput and MRPC accuracy of different fine-tuning methods. Token-level throughput (Thr., in k tokens/sec) is measured on LLaMA-2-7B using the GSM8K dataset with one A800 GPU, while MRPC classification accuracy (Acc., %) is evaluated on RoBERTa-large. Accuracy results are reported from the Block-Circulant Adapter (BCA) work [circulant], where only limited configurations were provided, leading to some “N/A” entries. All lora experiments use rank $r=32$. For circulant-based methods, p denotes the block size.